# Quantum computational speedup and retrocausality

### v3

@@ 2 Addenda
Before moving on to the formalizations of Section 3., let us open some of the “we will see” put for compactness in the Introduction.
@@
The initial measurement of the problem-setting in a quantum superposition (or mixture) of all the possible problem settings selects one of them at random; this is followed by the unitary computation of the corresponding solution (which never changes the problem-setting selected by the initial measurement) and the final measurement of the solution, which gives it with certainty. Of course, the problem-setting and the corresponding solution are two commuting observables. Note that the process that connects the initial with the final measurement outcome is unitary and thus time-reversible. As already mentioned, for shortness, we “confuse” the final measurement outcome with the state immediately before it (which coincides with it).
@@
The basic assumption is that the determination of the information that specifies the problem setting and the corresponding solution (one an invertible function of the other) evenly shares between the initial measurement of the problem setting and the final measurement of the solution, in a quantum superposition of all the possible ways of evenly sharing. In each time-symmetrization instance, the selection performed by the initial measurement propagates forward in time, that performed by the final measurement backward in time.
@@
As we will also see in the formalizations of Subsection 3.1, the time-symmetrization we are dealing with completes the ordinary (causal) quantum description of the quantum algorithm by:
@@
In the first place, let us note that the causal loops logically implied by the sheer existence of the quantum speedup, which tell the problem-solver half of the information about the solution she will produce and measure in the future, are, of course, with respect to her.
@@
In fact, with the kind of process we are dealing with, to describe a causal loop (any one), we must properly share the selection of the problem-setting between initial and final measurements, with the part of the selection performed by the initial measurement propagating forward in time and that performed by the final measurement backward in time. To obtain exactly the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup, this sharing must be even.
@@ 3 Formalization
This section is divided into two parts. In the first, we formalize the notion of completing the ordinary quantum description of quantum algorithm by time-symmetrizing it. The second is the generalization from quantum algorithms to quantum correlation.
@@ 3.1 Completing the ordinary quantum description of quantum algorithms by time-symmetrizing it
The completion of the quantum description of quantum algorithms consists of the following steps:
@@ 3.2 Generalization
First, everything immediately generalizes to any number, $N$, of drawers: it suffices to increase from $2=\sqrt{4}$ to $\sqrt{N}$ the number of quantum bits in each quantum register. We can see that all considerations remain unaltered: the problem-solver always knows in advance one of the possible halves of the information about the solution and can use this knowledge to produce the solution with correspondingly fewer computation steps, $\operatorname*{O}\left(\sqrt{N}\right)$ in the optimal logical and quantum case. In equivalent terms, the dimension of the Hilbert space in which the search for the solution – or for one-to-one correlation between the two observables – takes place reduces from $N$ basis vectors to $\sqrt{N}$ basis vectors.
@@ 4 Discussion
Starting with the already mentioned 1949 work of Wheeler and Feynman, various other works attempted and are attempting to go beyond the causal principle (i.e. one-way causality). However, we would think that their ways of going beyond have nothing to do with the time-symmetric causal loops of the present work. In our opinion, the discovery of the quantum speedup opens up an entirely new ground for the study of retrocausality.
@@
Opening to retrocausality, as in the present theory, can have, of course, a broad impact of many aspects of science. Of particular interested is its possible impact on the natural sciences, particularly, on the notion of teleological evolution.
@@
The point is that the present retrocausal explanation of the quantum speedup makes of quantum algorithms full fledged teleological evolutions, indeed evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
@@
As is well known, the teleological notion has been at the center of a long-standing debate between evolutionary biologists, physicists and natural philosophers. The question, of course, was whether the evolutions of the living are teleological in character. Eventually, the natural sciences dismissed the notion of teleological evolution right because it implies retrocausality, deemed to be a physical impossibility (not so natural philosophy, see further below).
Now, the fact quantum algorithms are full fledged teleological evolutions should already shake this position of the natural sciences. Moreover, the fact that the evolutions of the living belong to a macroscopic world whose classical description would obviously prevent their quantum speedup and consequent teleological character is not as definitive as it might seem.
This is because, under the quantum cosmological assumption, the physical description of the macroscopic reality would no more be one of classical physics.
The quantum cosmological assumption is the assumption that the universe as a whole, being a perfectly isolated system since by definition there is nothing outside it, is quantum and evolves in a unitary quantum way.
@@
In particular a quantum evolution of the universe, thus also its evolution toward life, would obviously involve quantum superpositions, and these, under the present theory, would imply retrocausality. An immediate, and trivial, result would therefore be the falsification of the statement that the evolution of the universe toward life cannot be teleological since this would require the physically impossible notion of retrocausality. Indeed, under the quantum cosmological assumption, and the present theory, the evolution of the universe toward life should host retrocausality.
@@
In $\left[22\right]$, we showed that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the evolution of the universe toward the “value of life” today (either “absence of life” or “existence of sentient life”) under the fundamental physical laws and their fundamental constants, would acquire the form of a quantum algorithm. Indeed, it would unitarily build up a (assume one-to-one) quantum correlation between the value of the fundamental constant since the beginning of time and the value of life today. We have seen that this is what a quantum algorithm does: indeed, it unitarily builds up a one-to-one correlation between the value of the problem-setting since its beginning and the value of the solution at its end. Either one could occur with quantum speedup and consequently be teleological in character.
By the way, in the quantum cosmological view, the description of the evolution of any macroscopic object could be obtained in principle from the quantum description of the evolution of the universe by tracing out all the rest. On the basis of this, always in $[22]$, we have shown that, in the universe we are in, the evolution of life on Earth and, with it, the evolutions of species, would inherit the teleological character of the evolution of the universe toward life.
@@
Of course, all the above is speculation, indeed like the quantum cosmological assumption. However, it would be enough to falsify the statement that teleological natural evolutions cannot have a physical basis. By the way, if the present argument in favor of the teleological character of the evolutions of the living goes against the current position of the natural sciences, instead it goes along with the position of natural philosophy, see $\left[25\right]$. It would be a case in which common sense prevailed over what, according to the present theory, would have been an unjustified precept of physics.
@@ 5 Conclusions
The basic assumption of the present work is that quantum algorithms are subject to classical logic. As a consequence, their ordinary quantum description could be flanked by a classical logic description, whereas, of course, the two descriptions should go along.
The most significant consequence of this assumption is that the classical logic description of a quantum state superposition, the latter both at the basis of the quantum speedup and one of the pillars of quantum mechanics, necessarily involves retrocausality. Let us recall how simple this is. Apparently, seeing the monk on the two banks of a river at the same time is a logical impossibility as the two possibilities are mutually exclusive with one another. However, it is an impossibility that vanishes by resorting to retrocausality. The monk, after reaching one bank, must have gone sufficiently backward in time undoing the path just followed and then forward in time again but on a different path that leads him to the other bank.
Consequent findings are: the retrocausal explanation of both the quantum speedup and quantum nolocality, the consequent teleological character of quantum algorithms and, under the (speculative) quantum cosmological assumption and in the universe we are in, the evolution of the universe toward life. The teleological character of the evolution of life on Earth and the evolutions of species would be consequent too.
Of course, this would be a significant change in a long standing view of physical reality. We would think that it is plausible enough to deserve a discussion.

### v4

@@ 2 Addenda
Before moving on to the formalizations of Section 3., let us open some of the “we will see” put for compactness in the previous section.
@@
The initial measurement of the problem-setting in a quantum superposition (or mixture) of all the possible problem settings selects one of them at random; this is followed by the unitary computation of the corresponding solution (which never changes the problem-setting selected by the initial measurement) and the final measurement of the latter. The solution is an invertible function of the problem-setting and is determined with cetainty.
Note that the process that connects the initial with the final measurement outcome is unitary and thus time-reversible. As already mentioned, for shortness, we “confuse” the final measurement outcome with the state immediately before it (which coincides with it).
@@
The basic assumption is that the selection of the problem setting and the corresponding solution (one an invertible function of the other) evenly shares between the initial measurement of the problem setting and the final measurement of the solution, in a quantum superposition of all the possible ways of evenly sharing. In each time-symmetrization instance, the selection performed by the initial measurement propagates forward in time, that performed by the final measurement backward in time.
@@
Summing up, the time-symmetrization we are dealing with completes the ordinary (causal) quantum description of the quantum algorithm by:
@@
In the first place, the causal loops logically implied by the sheer existence of the quantum speedup, which tell the problem-solver half of the information about the solution, are, of course, with respect to her.
@@
In fact, with the kind of process we are dealing with, to describe a causal loop (any one), we must properly share the selection of the problem-setting between initial and final measurements, with the part of the selection performed by the initial measurement propagating forward in time by the unitary computation of the solution and that performed by the final measurement backward in time by the inverse of that unitary computation. To obtain exactly the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup, this sharing must be even.
@@ 3 Formalization
This section is divided into two parts. In the first, we formalize the notion of completing the ordinary quantum description of the quantum algorithm by time-symmetrizing it. The second is the generalization from the quantum algorithm to quantum correlations.
@@ 3.1 Completing the ordinary quantum description of quantum algorithms by time-symmetrizing it
The completion of the quantum description of the quantum algorithm consists of the following steps:
@@ 3.2 Generalization
First, everything immediately generalizes to any number, $N$, of drawers: it suffices to increase from $2=\sqrt{4}$ to $\sqrt{N}$ the number of quantum bits in each quantum register. We can see that all considerations remain unaltered: the problem-solver always knows in advance one of the possible halves of the information about the solution and can use this knowledge to produce the solution with correspondingly fewer computation steps, $\operatorname*{O}\left(\sqrt{N}\right)$ in the optimal logical and quantum case. In equivalent terms, the dimension of the Hilbert space in which the search for the solution – or one-to-one correlation between the two observables – takes place reduces from $N$ basis vectors to $\sqrt{N}$ basis vectors.
@@ 4 Discussion
We mention the precedents to the present work and discuss possible further developments.
4.1 Precedents
Starting with the already mentioned 1949 Wheeler and Feynman’s work, various other works attempted and are attempting to go beyond the causal principle (i.e. one-way causality). However, we would think that their ways of going beyond have nothing to do with the time-symmetric causal loops of the present work. In our opinion, the discovery of the quantum speedup opens up an entirely new ground for the study of retrocausality.
@@ 4.1 Precedents
4.2 Looking forward
The form of retrocausality we are dealing with, which would be commonplace in ordinary quantum mechanics, would consequently have a broad impact on many aspects of science. Of special interest is that on the natural sciences, particularly on the notion of teleological evolution.
@@ 4.2 Looking forward
This is because the present retrocausal explanation of the quantum speedup makes of quantum algorithms full fledged teleological evolutions, indeed evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
@@
As is well known, the teleological notion has been the focus of a major debate between evolutionary biologists, physicists and natural philosophers. The question, of course, was whether the evolutions of the living can be teleological in character. Eventually, the natural sciences dismissed the notion of teleological evolution right because it implies a retrocausality deemed to be physically impossible (not so natural philosophy, see further below). That was not an easy decision. As is well known, the teleological notion perfectly matches with the functional structure of the living, for example “teeth were created for the purpose of chewing” (Aristotle). This difficulty is witnessed by the following sentence by the evolutionary biologist Haldane $\left[22\right]$:
Teleology is like a mistress to a biologist: he cannot live without her, but he is unwilling to be seen with her in public.
Now, the fact quantum algorithms are full fledged teleological evolutions should already shake this “official” position of the natural sciences. Furthermore, the fact the evolutions of the living belong to the macroscopic world described by a classical physics that cannot host any quantum speedup is not as definitive as it might seem. This is because, under the quantum cosmological assumption $\left[23,24\right]$, the classical physics description of the macroscopic reality is completely avoided. This assumption is that the universe as a whole, being a perfectly isolated system since by definition there is nothing outside it, is quantum and evolves in a unitary quantum way. As a consequence, in principle, the evolution of any (also macroscopic) system can be formally obtained from that of the universe by tracing out all the rest.
@@
In particular, a quantum evolution of the universe would obviously involve quantum superpositions, and these, under the present theory, imply retrocausality. An immediate consequence would be the falsification of the statement that the evolution of the universe toward life cannot be teleological since this would imply a physically impossible retrocausality. Indeed, under both the quantum cosmological assumption and the present theory, the evolution of the universe toward life should host retrocausality.
Moreover, always under the quantum cosmological assumption, the teleological character of quantum algorithms becomes a strong evidence of the possibility that the teleological evolutions of life are teleological in character $\left[25\right]$, as follows.
@@
Indeed, for the Fine-tuned Universe version of the Anthropic Principle, the evolution of the universe toward the “value of life” today (either “absence of life” or “existence of sentient life”) under the fundamental physical laws and their fundamental constants would acquire the form of a quantum algorithm. In fact, it would unitarily build up a (assume one-to-one) quantum correlation between the value of the fundamental constant since the beginning of time and the value of life today. This is what the quantum algorithms we have been dealing with777 With the solution and invertible function of the problem-setting and determined with certainty. do: they unitarily build up a one-to-one correlation between the value of the problem-setting since their beginning and the value of the solution at their end. Of course, this building up can always occur with quantum speedup and consequently be teleological in character. As shown in $\left[26\right]$, even a random quantum search for one-to-one correlation always occurs with quantum speedup.
In the universe we are in, what could be teleological is, of course, the evolution of the universe toward life. As a consequence – see $\left[25\right]$ – it could be teleological also the evolution of life on Earth (with it, the evolutions of species), the trace over all the rest of the teleological evolution of the universe toward life.
@@
Naturally, like the quantum cosmological assumption, all the above is speculation. However, being physically plausible, it should be enough to falsify the statement that teleological natural evolutions cannot have a physical basis. Note that, if this goes against the present “official” position of the natural sciences on the teleological issue, it goes instead perfectly along the present position of natural philosophy, see $\left[27\right]$. It would be a case where common sense prevailed over what, according to the present work, would turn out to have been a bias of science.
@@ 5 Conclusions
The starting point of the present work is that, involving only the measurements of commuting observables – the problem-setting and the corresponding solution – quantum algorithms should be subject to classical logic. Therefore, their ordinary quantum description could be flanked by a classical logic description and, of course, the two descriptions should go along.
We have seen that a classical logic description of a quantum state superposition – an essential part of quantum algorithms – necessarily involves a logical form of retrocausality that, in turn, must be physically implicit in it.
Let us recall the reason for this. Apparently, seeing the monk on the two banks of a river at the same time is a classical logic impossibility as, logically, the two possibilities are mutually exclusive with one another. However, it is an impossibility that vanishes by resorting to retrocausality. After reaching one bank, the monk must have gone sufficiently backward in time undoing the path just followed and then forward in time again but on a different path that led him to the other bank. This identically holds in the quantum world.
Summing up, the assumption that quantum algorithms should be describable in a classical logic way brings retrocausality into ordinary quantum mechanics, where it would be as common as quantum state superpositions. The form of retrocausality we are dealing with would correspondingly be pervasive:
It allows to provide a unified quantitative explanation of all quantum speedups and, together, quantum nonlocality.
It shows that the building up of one-to-one quantum correlation between two initially unrelated observables that does not change the value of the observable selected by the initial measurement can always occur with quadratic quantum speedup.
It highlights the teleological character of quantum algorithms and, under the quantum cosmological assumption, provides a plausible physical basis to the teleological character of natural evolutions.
Starting with the famous 1949 one by Wheeler and Feynman, various attempts have been done in the past to introduce the notion of retrocausality in physics. Of course, none of them ever entered mainstream physics. We would believe that the quantum computational speedup, a relatively recent discovery, opens up an entirely new ground for the study of retrocausality. Being biased, we would think that the universal character of the retrocausality found to be implicit in the quantum computational speedup and the (apparent) evidence of its existence make it worthy of discussion.